OpenStudy (anonymous):

4/x+12/x-5=-11 solve.

OpenStudy (anonymous):

i multiplied by the lcd and i got the answer 5 but it was wrong not sure what im doing wrong

OpenStudy (anonymous):

i multiplied by the lcd and i got the answer 5 but it was wrong not sure what im doing wrong

OpenStudy (anonymous):

16/x=-6 x=-16/6=-8/3

OpenStudy (anonymous):

\[\frac{4}{x}+\frac{12}{x}-5=-11\text{ }\text{ or }\text{ }\frac{ 4}{ x}+\frac{12}{x-5}=-11 \]

OpenStudy (anonymous):

(4/x)+(12/x-5) =-11

OpenStudy (anonymous):

(4/x)+(12/x-5) =-11

OpenStudy (anonymous):

(4/x)+(12/x-5) =-11

OpenStudy (anonymous):

In either case put all of the terms on the left hand side of the equation and combine the terms into one fraction. the fraction is zero therefore the numerator has to be zero. Set the numerator to = 0 and solve for x.

OpenStudy (anonymous):

that explanation didnt do anything for me

OpenStudy (anonymous):

(4/x) + (12/(x - 5)) = -11 Is the above the problem statement? This one is different than your last posting.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. If you bring the -11 over to the left hand side you will have: \[11+\frac{12}{-5+x}+\frac{4}{x}=0 \] The terms may be in a different order than yours, but the equation is still valid.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now when you combine all of the terms on the left hand side together, you will have the following fraction equal to zero: \[\frac{-20-39 x+11 x^2}{(-5+x) x}=0 \]

OpenStudy (anonymous):

???????????????

OpenStudy (anonymous):

???????????????

OpenStudy (anonymous):

??????? ... I get it. You are having trouble with combining fractions. Is that right?

OpenStudy (anonymous):

i dont see how you got the numerator

OpenStudy (anonymous):

4(x-5)+12(x)+11x(x-5)=0 -20-39x+11x^2=0 x=4 (+/_)

OpenStudy (anonymous):

my bad its only 4 next answer is -0.45

OpenStudy (anonymous):

You have to find a common divisor, an expression divisible by all of the denominators. A common divisor of 1, x-5 and x is x*(x-5). You then multiply each term on the left hand side by: The equation editor has crapped out at least on my version of Firefox. (x (x - 5))/(x (x - 5)) Multiply the above, which is equal to one, by each term on the right hand side. I have go to break off right now. Google "how to combine fractions algebra"

OpenStudy (anonymous):

Sorry. multiply each term on the left hand side. The solutions are x = -5/11 and x = 4 Got to go.